OpenStudy (anonymous):

If the hypotenuse of a 45º-45º-90º triangle is 54 inches long, what is the length of one of the legs? Leave your answer in radical form. 27 square root of 2 inches 54 square root of 2 Square root of 2 27 and can you tell me how to do this

OpenStudy (jamesj):

This is a right angled triangle and we can use Pythagorus' Theorem. And more than that, this is an isoceles triangle, so the two non-hypothenuse side lengths are the same. Call them X So X^2 + X^2 = 54^2 Now solve for X

OpenStudy (anonymous):

27?

OpenStudy (jamesj):

Let's test. If X = 27, then X^2 + X^2 = 2 . 27^2 But that is not equal to 54^2, because 54^2 = 2^2 . 27^2

OpenStudy (jamesj):

So no, 27 is wrong. Do the algebra and don't guess.

OpenStudy (anonymous):

27sqrt2?

OpenStudy (jamesj):

What did the algebra tell you?

OpenStudy (jamesj):

\[X^2 = (27\sqrt{2})^2 = 2 \times 27^2\] so\[X^2 + X^2 = 4 \times 27^2 = 2^2 \times 27^2 = (2 \times 27)^2 = 54^2\] Hence yes, 27.sqrt(2) is correct.